OpenStudy (luigi0210):

You are to design an open-top rectangular stainless-steel vat. It is to have a square base and a volume of 32ft^3. What dimensions would have the least material needed?

Elsa213 (elsa213):

x = side of square y = depth volume of open vat, V = x^2 y = 32 -------------- (1) to qualify the statement "to weigh no more than necessary means that the surface area (S) of plate = minimum S = x^2 + 4xy using (1) S = x^2 + 4x (32/x^2) = x^2 + 128/x dS/dx = 2x - 128/x^2 = 0 for minimum 2x^3 = 128 x^3 = 64 = 4^3 x = 4 ft y = 32/16 = 2 ft >>>>>>>> from (1) the dimensions are x = 4 ft of base side y = 2 ft as depth

Elsa213 (elsa213):

;-; i didnt even get medal <\3

TheSmartOne (thesmartone):

@Elsa213 Now you got a medal :P

Elsa213 (elsa213):

:D tank chu (~;u;~)

TheSmartOne (thesmartone):

:) So at least there has been 2 people who have stalked @Luigi0210 's question ;P